OpenStudy (anonymous):

The lengths of the sides of a rectangle become 4 times greater than the original measurements. Which statement is true about the rectangle? Its area becomes 4 times greater, and perimeter becomes 20 times greater. Its area becomes 16 times greater, and perimeter becomes 4 times greater. Its area becomes 4 times greater, and perimeter becomes 16 times greater. Its area becomes 20 times greater, and perimeter becomes 4 times greater

OpenStudy (anonymous):

The new area = 4l*4b=16(l*b) The new perimeter = 2(4l+4b)=4[2(l+b)] Its area becomes 16 times greater, and perimeter becomes 4 times greater.

OpenStudy (anonymous):

thnk you

OpenStudy (anonymous):

You're welcome.